Mapalad Integrated High School determined students’ Body Mass Index (BMI) at the opening of classes. It has been recorded that the average height of female students is 154.2 centimeters with a standard deviation of 9 centimeters. The researcher conducted her own study and she randomly selected 40 female students. In her study, she got an average of 156.7 centimeters. Is there a reason to believe the claims of the school? Use 5% level of significance in testing the hypothesis.



1
Expert's answer
2022-05-24T11:56:39-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=154.2"

"H_1:\mu\not=154.2"

This corresponds to a two-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha = 0.05," and the critical value for a two-tailed test is "z_c = 1.96."

The rejection region for this two-tailed test is "R = \{z:|z|>1.96\}."

The z-statistic is computed as follows:


"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\dfrac{156.7-154.2}{9/\sqrt{40}}\approx1.7568"

Since it is observed that "|z|=1.7568<1.96=z_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value is "p=2P(z>1.7568)=0.078952," and since "p= 0.078952>0.05=\alpha," it is concluded that the null hypothesis is notrejected.

Therefore, there is not enough evidence to claim that the population mean "\mu"

is different than 154.2, at the "\alpha = 0.05" significance level.